# Policy Optimization in the Linear Quadratic Gaussian Problem: A Frequency Domain Perspective

### v3

@@ 1 Introduction
Notably, the optimal dynamic filtering problem, as a simplified variant of LQG problem, gives rise to policy optimization methods with global convergence and sublinear convergence rates. Building on this problem, Umenberger et al. (2022) [32] propose the Differential Convex Lifting (DCL) method, which extends earlier convex parameterization theory. Their approach embeds the original optimization variables into an extended space and constructs a diffeomorphic mapping between the extended problem and a convex formulation, guaranteeing that all the stationary points in the extended domain correspond to globally optimal solutions of optimal dynamic filtering problem. Building on this theory, they prove that the optimal dynamic filtering process satisfies the weak Polyak-Lojasiewicz (PL) condition and further propose a transformation technique to ensure the coerciveness of the cost functional. The main theoretical tools for this work are drawn from the broader field of robust optimal control. In this line of research, Scherer et al. [28] develop convex parameterization methods for LQG problem, which form the basis of DCL. However, the effectiveness of these approaches is often restricted by their heavy reliance on accurate system models, which limits their applicability in structured $\mathcal{H}_{\infty}$ synthesis [2] and risk-sensitive scenarios [36]. These limitations, in turn, have motivated a recent surge of research on data-driven policy optimization methods. As a result, to address nonconvex problems such as LQG and dynamic filtering, researchers have begun to integrate policy optimization methods with convex parameterization theory to leverage the strengths of both methodologies. Zheng et al. (2023,2024) [39, 40] subsequently generalize DCL to LQG and $\mathcal{H}_{\infty}$ controller optimization. Their work demonstrates that the LQG problem is equivalent to a linearly constrained convex optimization problem. However, the construction of this problem requires accurate system models and assumptions of invertibility for model parameters.
@@ 3.1 The properties of problem (8)
$\mathcal{A}^{\top}\mathcal{P}_{\mathcal{K}}+\mathcal{P}_{\mathcal{K}}\mathcal{A}$ $=-\begin{bmatrix}Q&0\\ 0&C_{\mathcal{K}}^{\top}RC_{\mathcal{K}}\end{bmatrix}$,
$\mathcal{A}\Sigma_{\mathcal{K}}+\Sigma_{\mathcal{K}}\mathcal{A}^{\top}$ $=-\begin{bmatrix}\mathcal{W}&0\\ 0&B_{\mathcal{K}}\mathcal{V}B_{\mathcal{K}}^{\top}\end{bmatrix}$, (12)
@@ 3.2 Optimal condition analysis
[42] Let strictly proper real-coefficient rational vector functions $f(s)$ and $g(s)$ be square-integrable with no poles on the purely imaginary axis. Then, the $\mathcal{H}_{2}$ inner product $\langle f,g\rangle_{\mathcal{H}_{2}}$ equals the sum of residues of $f(-s)^{\top}g(s)$ at poles in the right half-plane: $Res_{\Re(s)\leq 0}\big(f(-s)^{\top}g(s)\big)$.
@@
$=\,\operatorname{tr}\left(\operatorname{Res}_{\Re(s)\leq 0}\Big(((\mathbf{M}_{12}^{\mathcal{K}}(-s))^{\top}\mathbf{M}_{11}^{\mathcal{K}}(s)(\mathbf{M}_{21}^{\mathcal{K}}(-s))^{\top}(\Delta\mathcal{K}(I-\mathbf{M}_{22}^{\mathcal{K}}(-s)\Delta\mathcal{K})^{-1})^{\top}\Big)\right).$
@@
$\mathbf{M}_{11}^{\mathcal{K}}(s)(\mathbf{M}_{21}^{\mathcal{K}}(-s))^{\top}=\left[\begin{array}[]{cc|c}\mathcal{A}&\mathcal{B}\mathcal{B}^{\top}&\mathcal{B}_{0}\\ 0&-\mathcal{A}^{\top}&\mathcal{C}_{0}\\ \hline\cr-\mathcal{C}&0&0\end{array}\right].$
@@
From Lyapunov equation (12), it follows that $\Sigma_{\mathcal{K}}\mathcal{A}^{\top}+\mathcal{A}\Sigma_{\mathcal{K}}+\mathcal{B}\mathcal{B}^{\top}=0$. By applying a similar similarity transformation to (LABEL:eq_real)
@@
$=\operatorname{tr}\big(Res_{\Re(s)\leq 0}\left(\mathcal{B}_{1}(sI+\mathcal{A}^{\top})^{-1}\mathcal{P}_{\mathcal{K}}(\mathcal{B}_{0}-\Sigma_{\mathcal{K}}\mathcal{C}_{0})(\Delta\mathcal{K}(I-\mathbf{M}_{22}^{\mathcal{K}}(-s)\Delta\mathcal{K})^{-1})^{\top}\right)$
$\hphantom{=}+Res_{\Re(s)\leq 0}\left((\mathcal{C}_{1}-\mathcal{B}_{1}\mathcal{P}_{\mathcal{K}})(sI-\mathcal{A})^{-1}(\mathcal{B}_{0}-\Sigma_{\mathcal{K}}\mathcal{C}_{0})(\Delta\mathcal{K}(I-\mathbf{M}_{22}^{\mathcal{K}}(-s)\Delta\mathcal{K})^{-1})^{\top}\right)$
$\hphantom{=}+Res_{\Re(s)\leq 0}\left(\mathcal{C}_{1}\Sigma_{\mathcal{K}}(sI+\mathcal{A}^{\top})^{-1}\mathcal{C}_{0}(\Delta\mathcal{K}(I-\mathbf{M}_{22}^{\mathcal{K}}(-s)\Delta\mathcal{K})^{-1})^{\top}\right)$
$\hphantom{=}+Res_{\Re(s)\leq 0}\left(\mathcal{B}_{1}(sI+\mathcal{A}^{\top})^{-1}\mathcal{C}^{\top}\mathcal{C}\Sigma_{\mathcal{K}}(sI+\mathcal{A}^{\top})^{-1}\mathcal{C}_{0}(\Delta\mathcal{K}(I-\mathbf{M}_{22}^{\mathcal{K}}(-s)\Delta\mathcal{K})^{-1})^{\top}\right)$
$=\operatorname{tr}\left(Res_{\Re(s)\leq 0}\left((\mathcal{C}_{1}-\mathcal{B}_{1}\mathcal{P}_{\mathcal{K}})(sI-\mathcal{A})^{-1}(\mathcal{B}_{0}-\Sigma_{\mathcal{K}}\mathcal{C}_{0})(\Delta\mathcal{K}(I-\mathbf{M}_{22}^{\mathcal{K}}(-s)\Delta\mathcal{K})^{-1})^{\top}\right)\right).$
Therefore, a sufficient condition for $(\mathbf{M}_{12}^{\mathcal{K}}(-s))^{\top}\mathbf{M}_{11}^{\mathcal{K}}(s)(\mathbf{M}_{21}^{\mathcal{K}}(-s))^{\top}$ to be anti-stable is that the second term in (3.2) vanishes. The preceding analysis has established the sufficiency of this condition.
@@
This condition ensures that the leading block of (27) vanishes. Consequently, it remains to analyze the remaining terms. Assume that $\omega_{0}\in\mathbb{R}$ makes $(\mathcal{C}_{1}-\mathcal{B}_{1}\mathcal{P}_{\mathcal{K}})(j\omega_{0}I-\mathcal{A})^{-1}(\mathcal{B}_{0}-\Sigma_{\mathcal{K}}\mathcal{C}_{0})\neq 0$. Then, we can consider a stable transfer function $\hat{\mathcal{F}}(s)$
@@
    The rank condition $\operatorname{rank}\left(\begin{bmatrix}P_{21}^{\mathcal{K}}&P_{22}^{\mathcal{K}}\end{bmatrix}\right)=\operatorname{rank}\left(\begin{bmatrix}\Sigma_{21}^{\mathcal{K}}&\Sigma_{22}^{\mathcal{K}}\end{bmatrix}\right)=q$ holds.
@@
$~\;=\begin{bmatrix}P_{1}^{\mathcal{K}}\mathcal{A}\Sigma_{2}^{\mathcal{K}\top}\\ P_{2}^{\mathcal{K}}\mathcal{A}\Sigma_{2}^{\mathcal{K}\top}\end{bmatrix}+\begin{bmatrix}A^{\top}_{\mathcal{K}}P_{1}^{\mathcal{K}}\Sigma_{2}^{\mathcal{K}\top}\\ C_{\mathcal{K}}^{\top}B^{\top}P_{1}^{\mathcal{K}}\Sigma_{2}^{\mathcal{K}\top}\end{bmatrix}$
@@
$\begin{bmatrix}P_{1}^{\mathcal{K}}\mathcal{A}\Sigma_{2}^{\mathcal{K}\top}\\ P_{2}^{\mathcal{K}}\mathcal{A}\Sigma_{2}^{\mathcal{K}\top}\end{bmatrix}+\begin{bmatrix}A^{\top}_{\mathcal{K}}P_{1}^{\mathcal{K}}\Sigma_{2}^{\mathcal{K}\top}\\ C_{\mathcal{K}}^{\top}\frac{\nabla_{C_{\mathcal{K}}}J(\mathcal{K})}{2}\end{bmatrix}=\begin{bmatrix}P_{1}^{\mathcal{K}}\mathcal{A}\Sigma_{2}^{\mathcal{K}\top}\\ P_{2}^{\mathcal{K}}\mathcal{A}\Sigma_{2}^{\mathcal{K}\top}\end{bmatrix}+\begin{bmatrix}A^{\top}_{\mathcal{K}}P_{1}^{\mathcal{K}}\Sigma_{2}^{\mathcal{K}\top}\\ 0\end{bmatrix}=-\begin{bmatrix}Q\Sigma_{12}^{\mathcal{K}}\\ 0\end{bmatrix}$.
@@
To further exploit this structure, we construct a transformation matrix based on the rank condition $\operatorname{rank}(\Sigma_{2}^{\mathcal{K}})=\operatorname{rank}(P_{2}^{\mathcal{K}})=q$. Specifically, let
@@ 4 The difference of optimality conditions of LQG and LQR
where $\mathbb{K}_{LQR}$ denotes the set of all the state-feedback gains that stabilize system (LABEL:op_lqr), and $\{\omega_{t}^{\prime}\}$ is the Gaussian white noise process with intensity matrix $I$; the remaining notation is consistent with that in (4).
@@
To illustrate the explanatory power of our framework, we now use it to deconstruct the well-known difference between the LQR and LQG optimization landscapes. We begin with the LQR case. Unlike the general LQG problem, problem (LABEL:op_lqr) is known to admit a unique stationary point, which is also globally optimal [9]. The following theorem formalizes the insight gained from our framework (Theorem 3.4): it demonstrates that the structure of LQR problem is precisely what forces any stationary point to satisfy the global optimality condition. This perspective not only recovers the classic result but also sets the stage for our subsequent analysis in Remark 4.2, which will pinpoint exactly why suboptimal stationary points arise in the more challenging LQG setting.
@@
$=\left[\begin{array}[]{ccc|c}-(A+BK)^{\top}&Q+K^{\top}RK&0&0\\ 0&A+BK&I&0\\ 0&0&-(A+BK)^{\top}&I\\ \hline\cr B^{\top}&RK&0&0\end{array}\right].$
@@
$=\!\!\!\left[\begin{array}[]{ccc|c}-(A+BK)^{\top}&0&(Q+K^{\top}RK)\Sigma_{K}&P_{K}\Sigma_{K}\\ 0&(A+BK)&0&-\Sigma_{K}\\ 0&0&-(A+BK)^{\top}&I\\ \hline\cr B^{\top}&-BP_{K}+RK&RK\Sigma_{K}&0\end{array}\right],$
@@
$=\operatorname{tr}\left(\operatorname{Res}_{\Re(s)\leq 0}\left((-BP_{K}+RK)(sI-(A+BK))^{-1}\Sigma_{K}(\Delta K(I-\mathbf{M}_{22}^{K}(-s)\Delta K)^{-1})^{\top}\right)\right)$
$=\operatorname{tr}\left(\operatorname{Res}_{\Re(s)\leq 0}\left((-BP_{K}+RK)(sI-(A+BK))^{-1}\Sigma_{K}\Delta K^{\top}\right)\right)+o(\|\Delta K\|)$
@@
$(-BP_{K}+RK)(sI-(A+BK))^{-1}\Sigma_{K}\equiv 0,\quad\forall s\in\mathbb{C}.$ (57)
@@
$\nabla_{K}f_{LQR}(K)$ $=2\operatorname{Res}_{\Re(s)\leq 0}\left((-BP_{K}+RK)(sI-(A+BK))^{-1}\Sigma_{K}\right)$ (58)
$=2(-BP_{K}+RK)\Sigma_{K}.$
Equation (58) shows that the gradient of (LABEL:op_lqr) with respect to $K$ is precisely the residue associated with the optimality condition in (56). Since $\Sigma_{K}\succ 0$, the transfer function in (56) is absolutely controllable. Therefore, (56) (and hence the gradient) vanishes if and only if $(-BP_{K}+RK)\Sigma_{K}=0$, i.e., the system in (56) is absolutely unobservable. Consequently, all the stationary points of problem (LABEL:op_lqr) correspond to the optimal solutions. Moreover, the unique solution is characterized by $BP_{K}+RK=0$ together with the Lyapunov equation (52), which completes the proof. ∎
@@
$\nabla J(\mathcal{K})=2\mathbf{e}_{(q+m_{1})\times(q+m_{2})}\big(\operatorname{Res}_{\Re(s)\leq 0}((\mathcal{C}_{1}-\mathcal{B}_{1}\mathcal{P}_{\mathcal{K}})(sI-\mathcal{A})^{-1}(\mathcal{B}_{0}-\Sigma_{\mathcal{K}}\mathcal{C}_{0})),\,m_{1},m_{2}\big),$ (59)

### v4

@@ 1 Introduction
Notably, the optimal dynamic filtering problem, as a simplified variant of LQG problem, gives rise to policy optimization methods with global convergence and sublinear convergence rates. Building on this problem, Umenberger et al. (2022) [34] propose the Differential Convex Lifting (DCL) method, which extends earlier convex parameterization theory. Their approach embeds the original optimization variables into an extended space and constructs a diffeomorphic mapping between the extended problem and a convex formulation, guaranteeing that all the stationary points in the extended domain correspond to globally optimal solutions of optimal dynamic filtering problem. Building on this theory, they prove that the optimal dynamic filtering process satisfies the weak Polyak-Lojasiewicz (PL) condition and further propose a transformation technique to ensure the coerciveness of the cost functional. The main theoretical tools for this work are drawn from the broader field of robust optimal control. In this line of research, Scherer et al. [30] develop convex parameterization methods for LQG problem, which form the basis of DCL. However, the effectiveness of these approaches is often restricted by their heavy reliance on accurate system models, which limits their applicability in structured $\mathcal{H}_{\infty}$ synthesis [2] and risk-sensitive scenarios [38]. These limitations, in turn, have motivated a recent surge of research on data-driven policy optimization methods. As a result, to address nonconvex problems such as LQG and dynamic filtering, researchers have begun to integrate policy optimization methods with convex parameterization theory to leverage the strengths of both methodologies. Zheng et al. (2023,2024) [41, 42] subsequently generalize DCL to LQG and $\mathcal{H}_{\infty}$ controller optimization. Their work demonstrates that the LQG problem is equivalent to a convex optimization problem. However, the construction of this problem requires accurate system models and assumptions of invertibility for model parameters. Moreover, several works explore a novel parameterization method via input-output historical data for the discrete-time LQG problem [29, 7]. These approaches formulate an objective function that satisfies the PL condition, thereby guaranteeing linear convergence for policy gradient methods. Nevertheless, extending this framework to continuous-time systems has yet to be addressed.
@@ 3.1 The properties of problem (8)
$\mathcal{A}^{\top}\mathcal{P}_{\mathcal{K}}+\mathcal{P}_{\mathcal{K}}\mathcal{A}$ $=-\begin{bmatrix}Q&0\\ 0&C_{\mathcal{K}}^{\top}RC_{\mathcal{K}}\end{bmatrix}$, (12)
$\mathcal{A}\Sigma_{\mathcal{K}}+\Sigma_{\mathcal{K}}\mathcal{A}^{\top}$ $=-\begin{bmatrix}\mathcal{W}&0\\ 0&B_{\mathcal{K}}\mathcal{V}B_{\mathcal{K}}^{\top}\end{bmatrix}$,
@@ 3.2 Optimal condition analysis
[44] Let strictly proper real-coefficient rational vector functions $f(s)$ and $g(s)$ be square-integrable with no poles on the purely imaginary axis. Then, the $\mathcal{H}_{2}$ inner product $\langle f,g\rangle_{\mathcal{H}_{2}}$ equals the sum of residues of $f(-s)^{\top}g(s)$ at poles in the right half-plane: $\operatorname{Res}_{\Re(s)<0}\big(f(-s)^{\top}g(s)\big)$.
@@
$=\,\operatorname{tr}\left(\operatorname{Res}_{\Re(s)<0}\Big(((\mathbf{M}_{12}^{\mathcal{K}}(-s))^{\top}\mathbf{M}_{11}^{\mathcal{K}}(s)(\mathbf{M}_{21}^{\mathcal{K}}(-s))^{\top}(\Delta\mathcal{K}(I-\mathbf{M}_{22}^{\mathcal{K}}(-s)\Delta\mathcal{K})^{-1})^{\top}\Big)\right).$
@@
$\mathbf{M}_{11}^{\mathcal{K}}(s)(\mathbf{M}_{21}^{\mathcal{K}}(-s))^{\top}=\left[\begin{array}[]{cc|c}\mathcal{A}&\mathcal{B}\mathcal{B}^{\top}&\mathcal{B}_{0}\\ 0&-\mathcal{A}^{\top}&\mathcal{C}_{0}\\ \hline\cr\mathcal{C}&0&0\end{array}\right].$
@@
$(\mathbf{M}_{12}^{\mathcal{K}}(-s))^{\top}\mathbf{M}_{11}^{\mathcal{K}}(s)(\mathbf{M}_{21}^{\mathcal{K}}(-s))^{\top}$ (31)
@@
From Lyapunov equation (12), it follows that $\Sigma_{\mathcal{K}}\mathcal{A}^{\top}+\mathcal{A}\Sigma_{\mathcal{K}}+\mathcal{B}\mathcal{B}^{\top}=0$. By applying a similar similarity transformation to (31):
@@
$(\mathbf{M}_{12}^{\mathcal{K}}(-s))^{\top}\mathbf{M}_{11}^{\mathcal{K}}(s)(\mathbf{M}_{21}^{\mathcal{K}}(-s))^{\top}$ (33)
@@
$=\operatorname{tr}\big(\operatorname{Res}_{\Re(s)<0}\left(\mathcal{B}_{1}(sI+\mathcal{A}^{\top})^{-1}\mathcal{P}_{\mathcal{K}}(\mathcal{B}_{0}-\Sigma_{\mathcal{K}}\mathcal{C}_{0})(\Delta\mathcal{K}(I-\mathbf{M}_{22}^{\mathcal{K}}(-s)\Delta\mathcal{K})^{-1})^{\top}\right)$
$\hphantom{=}+\operatorname{Res}_{\Re(s)<0}\left((\mathcal{C}_{1}-\mathcal{B}_{1}\mathcal{P}_{\mathcal{K}})(sI-\mathcal{A})^{-1}(\mathcal{B}_{0}-\Sigma_{\mathcal{K}}\mathcal{C}_{0})(\Delta\mathcal{K}(I-\mathbf{M}_{22}^{\mathcal{K}}(-s)\Delta\mathcal{K})^{-1})^{\top}\right)$
$\hphantom{=}+\operatorname{Res}_{\Re(s)<0}\left(\mathcal{C}_{1}\Sigma_{\mathcal{K}}(sI+\mathcal{A}^{\top})^{-1}\mathcal{C}_{0}(\Delta\mathcal{K}(I-\mathbf{M}_{22}^{\mathcal{K}}(-s)\Delta\mathcal{K})^{-1})^{\top}\right)$
$\hphantom{=}+\operatorname{Res}_{\Re(s)<0}\left(\mathcal{B}_{1}(sI+\mathcal{A}^{\top})^{-1}\mathcal{C}^{\top}\mathcal{C}\Sigma_{\mathcal{K}}(sI+\mathcal{A}^{\top})^{-1}\mathcal{C}_{0}(\Delta\mathcal{K}(I-\mathbf{M}_{22}^{\mathcal{K}}(-s)\Delta\mathcal{K})^{-1})^{\top}\right)$
$=\operatorname{tr}\left(\operatorname{Res}_{\Re(s)<0}\left((\mathcal{C}_{1}-\mathcal{B}_{1}\mathcal{P}_{\mathcal{K}})(sI-\mathcal{A})^{-1}(\mathcal{B}_{0}-\Sigma_{\mathcal{K}}\mathcal{C}_{0})(\Delta\mathcal{K}(I-\mathbf{M}_{22}^{\mathcal{K}}(-s)\Delta\mathcal{K})^{-1})^{\top}\right)\right).$
Therefore, a sufficient condition for $(\mathbf{M}_{12}^{\mathcal{K}}(-s))^{\top}\mathbf{M}_{11}^{\mathcal{K}}(s)(\mathbf{M}_{21}^{\mathcal{K}}(-s))^{\top}$ to be anti-stable is that the second term in (33) vanishes. The preceding analysis has established the sufficiency of this condition.
@@
This condition ensures that the leading block of (27) vanishes. Consequently, it remains to analyze the remaining terms. We can consider a stable transfer function $\hat{\mathcal{F}}(s)$
@@
    The rank condition $\operatorname{rank}\left(\begin{bmatrix}P_{21}^{\mathcal{K}}&P_{22}^{\mathcal{K}}\end{bmatrix}\right)=\operatorname{rank}\left(\begin{bmatrix}\Sigma_{21}^{\mathcal{K}}&\Sigma_{22}^{\mathcal{K}}\end{bmatrix}\right)=q=n$ holds.
@@
$~\;=\begin{bmatrix}P_{1}^{\mathcal{K}}\mathcal{A}\Sigma_{2}^{\mathcal{K}\top}\\ P_{2}^{\mathcal{K}}\mathcal{A}\Sigma_{2}^{\mathcal{K}\top}\end{bmatrix}+\begin{bmatrix}A^{\top}P_{1}^{\mathcal{K}}\Sigma_{2}^{\mathcal{K}\top}\\ C_{\mathcal{K}}^{\top}B^{\top}P_{1}^{\mathcal{K}}\Sigma_{2}^{\mathcal{K}\top}\end{bmatrix}$
@@
$\begin{bmatrix}P_{1}^{\mathcal{K}}\mathcal{A}\Sigma_{2}^{\mathcal{K}\top}\\ P_{2}^{\mathcal{K}}\mathcal{A}\Sigma_{2}^{\mathcal{K}\top}\end{bmatrix}+\begin{bmatrix}A^{\top}P_{1}^{\mathcal{K}}\Sigma_{2}^{\mathcal{K}\top}\\ C_{\mathcal{K}}^{\top}\frac{\nabla_{C_{\mathcal{K}}}J(\mathcal{K})}{2}\end{bmatrix}=\begin{bmatrix}P_{1}^{\mathcal{K}}\mathcal{A}\Sigma_{2}^{\mathcal{K}\top}\\ P_{2}^{\mathcal{K}}\mathcal{A}\Sigma_{2}^{\mathcal{K}\top}\end{bmatrix}+\begin{bmatrix}A^{\top}P_{1}^{\mathcal{K}}\Sigma_{2}^{\mathcal{K}\top}\\ 0\end{bmatrix}=-\begin{bmatrix}Q\Sigma_{12}^{\mathcal{K}}\\ 0\end{bmatrix}$.
@@
To further exploit this structure, we construct a transformation matrix based on the rank condition $\operatorname{rank}(\Sigma_{2}^{\mathcal{K}})=\operatorname{rank}(P_{2}^{\mathcal{K}})=n$. Specifically, let
@@ 4 The difference of optimality conditions of LQG and LQR
where $\mathbb{K}_{LQR}$ denotes the set of all the state-feedback gains stabilizing system (LABEL:op_lqr), and $\{\omega_{t}^{\prime}\}$ is the Gaussian white noise process with intensity matrix $I$; the remaining notation is consistent with that in (4).
@@
To illustrate the explanatory power of our framework, we now use it to deconstruct the well-known difference between the LQR and LQG optimization landscapes. We begin with the LQR case. Unlike the general LQG problem, problem (LABEL:op_lqr) is known to admit a unique stationary point, which is also globally optimal [10]. The following theorem formalizes the insight gained from our framework (Theorem 3.4). Although this result is well known and has been previously established, by applying the new proof method developed through the analysis in Section 3, we demonstrate that the inherent structure of the LQR problem ensures that any stationary point satisfies the global optimality condition. This perspective not only recovers the classic result but also sets the stage for our subsequent analysis in Remark 4.2, which will pinpoint exactly why suboptimal stationary points arise in the more challenging LQG setting.
@@
$=\left[\begin{array}[]{ccc|c}-(A+BK)^{\top}&Q+K^{\top}RK&0&0\\ 0&A+BK&I&0\\ 0&0&-(A+BK)^{\top}&-I\\ \hline\cr-B^{\top}&RK&0&0\end{array}\right].$
@@
$=\!\!\!\left[\begin{array}[]{ccc|c}-(A+BK)^{\top}&0&(Q+K^{\top}RK)\Sigma_{K}&P_{K}\Sigma_{K}\\ 0&(A+BK)&0&\Sigma_{K}\\ 0&0&-(A+BK)^{\top}&-I\\ \hline\cr-B^{\top}&B^{\top}P_{K}+RK&RK\Sigma_{K}&0\end{array}\right],$
@@
$=\operatorname{tr}\left(\operatorname{Res}_{\Re(s)<0}\left((B^{\top}P_{K}+RK)(sI-(A+BK))^{-1}\Sigma_{K}(\Delta K(I-\mathbf{M}_{22}^{K}(-s)\Delta K)^{-1})^{\top}\right)\right)$
$=\operatorname{tr}\left(\operatorname{Res}_{\Re(s)<0}\left((B^{\top}P_{K}+RK)(sI-(A+BK))^{-1}\Sigma_{K}\Delta K^{\top}\right)\right)+o(\|\Delta K\|)$
@@
$(B^{\top}P_{K}+RK)(sI-(A+BK))^{-1}\Sigma_{K}\equiv 0,\quad\forall s\in\mathbb{C}.$ (59)
@@
$\nabla_{K}f_{LQR}(K)$ $=2\operatorname{Res}_{\Re(s)<0}\left((B^{\top}P_{K}+RK)(sI-(A+BK))^{-1}\Sigma_{K}\right)$ (60)
$=2(B^{\top}P_{K}+RK)\Sigma_{K}.$
Equation (60) shows that the gradient of (LABEL:op_lqr) with respect to $K$ is precisely the residue associated with the optimality condition in (58). Since $\Sigma_{K}\succ 0$, the transfer function in (58) is absolutely controllable. Therefore, (58) (and hence the gradient) vanishes if and only if $(B^{\top}P_{K}+RK)\Sigma_{K}=0$, i.e., the system in (58) is absolutely unobservable. Consequently, all the stationary points of problem (LABEL:op_lqr) correspond to the optimal solutions. Moreover, the unique solution is characterized by $B^{\top}P_{K}+RK=0$ together with the Lyapunov equation (54), which completes the proof. ∎
@@
$\nabla J(\mathcal{K})=2\mathbf{e}_{(q+m_{1})\times(q+m_{2})}\big(\operatorname{Res}_{\Re(s)<0}((\mathcal{C}_{1}-\mathcal{B}_{1}\mathcal{P}_{\mathcal{K}})(sI-\mathcal{A})^{-1}(\mathcal{B}_{0}-\Sigma_{\mathcal{K}}\mathcal{C}_{0})),\,m_{1},m_{2}\big),$ (61)